# An Advanced Deep Learning Framework for Ischemic and Hemorrhagic Brain Stroke Diagnosis Using Computed Tomography (CT) Images

### v1

@@ 3.5 Model Selection
4 Experimental Setup & Implementation
4.1 Environmental Setting
The experimental study was conducted using a Windows 10 operating system on a ×64 processor, intel(R) core i5 8th Gen CPU running at 1.60GHz to 2.71GHz with 8 GB of DDR4 RAM. Anaconda Navigator was used with the Jupyter Notebook for image processing applications.
@@ 4.3 Implementation Details
The proposed brain stroke diagnosis model begins with collecting and preparing datasets, Dataset-1 and Dataset-2, which are then manually curated to ensure data quality. This includes removing irrelevant and duplicate entries and incorporating additional data from local clinic to enhance the dataset’s diversity and representativeness. The preprocessing stage involves resizing each image to consistent dimensions of 224×224 pixels for most CNN architectures and 299×299 pixels for Inception V3 and Xception. Rescaling is performed to normalize the pixel values to a range of [0, 1], ensuring compatibility with the input requirements of the candidate models. Following preprocessing, feature extraction is conducted using five pre-trained convolutional neural network (CNN) architectures: DenseNet201, InceptionV3, Xception, ResNet50, and MobileNet V2. The classification layers of these CNNs are removed, allowing the penultimate layer’s feature maps to be extracted. These extracted features are saved as high-dimensional vectors in a CSV format for further processing. The feature vectors are then subjected to either feature selection and dimensionality reduction techniques or directly used without feature selection. Feature reduction methods include Bacterial Foraging Optimization (BFO), A bio-inspired optimization technique for selecting the most relevant features. Principal Component Analysis (PCA) is a statistical method to reduce dimensionality while preserving variance. Linear Discriminant Analysis (LDA) is a technique to maximize class separability by projecting data into a lower-dimensional space. The processed features are subsequently fed into a set of traditional machine learning classifiers: Support Vector Classifier (SVC), Random Forest (RF), Gaussian Naive Bayes (GNB), Decision Tree (DT), Extreme Gradient Boosting (XGB), K-Nearest Neighbors (KNN), and Logistic Regression (LR). Each classifier is trained and evaluated to determine its performance in predicting stroke categories. Finally, the best-performing model is selected based on its performance metrics and is utilized to classify outcomes into three categories: hemorrhagic stroke, ischemic stroke, and normal condition. This end-to-end approach ensures a robust and accurate diagnosis system suitable for real-world clinical applications. The following algorithm 1 shows the implementation procedure.
@@
4.4 Performance Evaluation Matrices
@@ 5.1 Result Analysis
5.2 Discussions & Limitations
In this study, we proposed a feature extraction-based approach utilizing pre-trained deep learning models combined with various classifiers and feature optimization techniques to enhance classification performance. Instead of relying on computationally expensive end-to-end deep learning models, we leveraged transfer learning to extract meaningful features from brain imaging data, which were then fed into traditional machine learning classifiers for stroke classification. Our extensive experimentation with multiple model-classifier combinations revealed that MobileNetV2+LDA+SVC was the best-performing model. This combination achieved the highest accuracy, precision, recall, and F1-score while maintaining a strong trade-off between computational efficiency and classification performance. Additionally, the model demonstrated robustness across different feature selection techniques, further validating its reliability for stroke detection. Compared to other deep learning models, MobileNetV2 is a lightweight architecture, making it a practical choice for real-world applications where computational resources may be limited.
We compared our findings with prior research on brain stroke detection using computed tomography (CT) scan images, our model outperformed most existing studies in this domain. While some studies, such as [20, 27, 30, 32] used Hybrid CBAM, EfficientNet-B0, AlexNet and UNet, reported slightly higher accuracy. it is important to note that these models likely employed end-to-end training strategies that require extensive computational resources and large-scale datasets for optimal performance. In contrast, our feature extraction-based method offers a more efficient and scalable solution, as it eliminates the need for training deep models from scratch. Additionally, many of the high-performing models in prior studies do not report computational complexity,
@@ 5.2 Discussions & Limitations
Table 10: Comparison with Prior Works on Brain Stroke Diagnosis
Ref. Year Model Findings
A. Abumihsan et al. [20] 2025 Hybrid CBAM 99.21%
M.M Hossain et al. [21] 2025 ViT-LSTM 96.61%
I. Tahyudin et al. [22] 2025 CBAM+ResNet18 95.00%
UmaMaheswaran et al. [23] 2024 XGB 97.00%
M.A. Saleem et al. [24] 2024 BiLSTM 96.50%
M. Sabir et al. [25] 2024 DCNN 96.50%
B. Borsos et al. [26] 2024 DAFT 80.00%
S. Prasher et al. [27] 2024 EfficientNet-B0 98.72%
C.D. Kulathilake et al. [28] 2024 DenseNet201 98.02%
Y. Ahmed et al. [29] 2024 3D CNN 92.50%
F. Yousaf et al. [30] 2023 UNet 99.57%
R. Raj et al. [31] 2023 StrokeViT 92.00%
A. Diker et al. [32] 2023 AlexNet 99.90%
[This Study] 2025 MobileNet+LDA+SVC 97.93%
@@
making it difficult to assess their feasibility for deployment in real-time clinical settings. By incorporating feature optimization techniques such as PCA, LDA, and BFO, our approach not only enhances classification accuracy but also reduces the risk of overfitting and improves generalization across different dataset sizes. Another key advantage of our model is its ability to balance high classification performance with computational efficiency, making it a viable option for deployment in medical imaging systems and resource-constrained environments.
@@ 6 Conclusion
This study explored the use of deep feature extraction combined with machine learning classifiers and optimization techniques to develop an improved stroke detection system. Through extensive experimentation, we identified MobileNetV2+SVC+LDA as the most effective model, achieving high classification accuracy of 97.93%, while maintaining computational efficiency. The proposed approach effectively balances performance, interpretability, and resource constraints, making it a suitable choice for stroke classification tasks. Our findings demonstrate the advantages of using feature extraction-based models over computationally expensive end-to-end deep learning frameworks. By leveraging Linear Discriminant Analysis (LDA) for dimensionality reduction and Support Vector Classifier (SVC) for classification, our system enhances both accuracy and inference speed. The results confirm that lightweight architectures like MobileNetV2, when paired with appropriate classifiers and optimization techniques, can deliver state-of-the-art performance without excessive computational demands. Future research could explore self-supervised learning and transformer-based architectures, which have shown promise in medical imaging tasks. Additionally, incorporating multi-modal data sources, such as clinical history and radiological findings, could improve diagnostic accuracy. Expanding the dataset with more diverse and real-world stroke cases will further validate the model’s generalizability. Finally, developing a real-time deployment system with interpretability features will enhance clinical adoption, aiding medical professionals in fast and accurate stroke diagnosis.

### v2

@@ 3.5 Classification Models
4 Experimental Setup and Implementation
4.1 Environment
The experimental study was conducted using a Windows 10 operating system on a ×64 processor, intel(R) core i5 8th Gen CPU running at 1.60GHz to 3.71GHz with 8 GB of DDR4 RAM. Anaconda Navigator was used with the Jupyter Notebook for image processing applications.
@@ 4.3 Implementation Details
The proposed brain stroke diagnosis model begins with collecting and preparing datasets, Dataset-1 and Dataset-2, which are then manually curated to ensure data quality. This includes removing irrelevant and duplicate entries and incorporating additional data from local clinic to enhance the dataset’s diversity
@@
and representativeness. The preprocessing stage involves resizing each image to consistent dimensions of 224×224 pixels for most CNN architectures and 299×299 pixels for Inception V3 and Xception. Rescaling is performed to normalize the pixel values to a range of [0, 1], ensuring compatibility with the input requirements of the candidate models. Following preprocessing, feature extraction is conducted using five pre-trained convolutional neural network (CNN) architectures: DenseNet201, InceptionV3, Xception, ResNet50, and MobileNet V2. The classification layers of these CNNs are removed, allowing the penultimate layer’s feature maps to be extracted. These extracted features are saved as high-dimensional vectors in a CSV format for further processing. The feature vectors are then subjected to either feature selection and dimensionality reduction techniques or directly used without feature selection. Feature reduction methods include Bacterial Foraging Optimization (BFO), A bio-inspired optimization technique for selecting the most relevant features. Principal Component Analysis (PCA) is a statistical method to reduce dimensionality while preserving variance. Linear Discriminant Analysis (LDA) is a technique to maximize class separability by projecting data into a lower-dimensional space. The processed features are subsequently fed into a set of traditional machine learning classifiers: Support Vector Classifier (SVC), Random Forest (RF), Gaussian Naive Bayes (GNB), Decision Tree (DT), Extreme Gradient Boosting (XGB), K-Nearest Neighbors (KNN), and Logistic Regression (LR). Each classifier is trained and evaluated to determine its performance in predicting stroke categories. Finally, the best-performing model is selected based on its performance metrics and is utilized to classify outcomes into three categories: hemorrhagic stroke, ischemic stroke, and normal condition. This end-to-end approach ensures a robust and accurate diagnosis system suitable for real-world clinical applications. The algorithm 1 shows the implementation procedure of the proposed brain stroke detection method.
4.4 Evaluation Matrices
@@ 5.1 Result Analysis
5.2 Discussion and Limitations
In this study, we proposed a feature extraction-based approach utilizing pre-trained deep learning models combined with various classifiers and feature optimization techniques to enhance classification performance.
@@ 5.2 Discussion and Limitations
Table 10: Comparison with State-of-the-Art Methods for Brain Stroke Diagnosis Using Computed Tomography Images
Ref. Dataset Class Accuracy Feature Optimization Cross validation Time&Space Complexity Computational Cost
A. Abumihsan et al. [20] 2000 & 2501 2 99.21% Moderate
M.M Hossain et al. [21] 770 & 2501 2 96.61% High
I. Tahyudin et al. [22] 2501 2 95.00% High
UmaMaheswaran et al. [23] 2501 2 97.00% Low
M.A. Saleem et al. [24] 2501 2 96.50% High
M. Sabir et al. [25] 10000 2 96.50% High
S. Prasher et al. [26] 2501 2 98.72% High
C.D. Kulathilake et al. [27] 4767 2 98.02% High
R. Raj et al. [28] 2431 2 92.00% High
A. Diker et al. [29] 2501 2 97.06% High
[This Study] 3819 3 97.93% Low
Instead of relying on computationally expensive end-to-end deep learning models, we leveraged transfer learning to extract meaningful features from brain imaging data, which were then fed into traditional machine learning classifiers for stroke classification. Our extensive experimentation with multiple model-classifier combinations revealed that MobileNetV2+LDA+SVC was the best-performing model. This combination achieved the highest accuracy, precision, recall, and F1-score while maintaining a strong trade-off between computational efficiency and classification performance. Additionally, the model demonstrated robustness across different feature selection techniques, further validating its reliability for stroke detection. Compared to other deep learning models, MobileNetV2 is a lightweight architecture, making it a practical choice for real-world applications where computational resources may be limited.
@@
We compared our findings with prior research on brain stroke detection using computed tomography (CT) images; our model outperformed most existing studies in this domain. While some studies, such as [20, 26, 27] used Hybrid CBAM, EfficientNet-B0, and DenseNet201, reported slightly higher accuracy. it is important to note that these models likely employed end-to-end training strategies that require extensive computational resources and large-scale datasets for optimal performance. In contrast, our feature extraction-based method offers a more efficient and scalable solution, as it eliminates the need for training deep models from scratch. Additionally, many of the high-performing models in prior studies do not report computational complexity, making it difficult to assess their feasibility for deployment in real-time clinical settings. By incorporating feature optimization techniques such as PCA, LDA, and BFO, our approach not only enhances classification accuracy but also reduces the risk of overfitting and improves generalization across different dataset sizes. Another key advantage of our model is its ability to balance high classification performance with computational efficiency, making it a viable option for deployment in medical imaging systems and resource-constrained environments.
@@ 6 Conclusion
This study explored the use of deep feature extraction combined with machine learning classifiers and optimization techniques to develop an improved stroke detection system. Through extensive experimentation, we identified MobileNetV2+SVC+LDA as the most effective model, achieving a high classification accuracy of 97.93%, while maintaining computational efficiency. The proposed approach effectively balances performance, interpretability, and resource constraints, making it a suitable choice for stroke classification tasks. Our findings demonstrate the advantages of using feature extraction-based models over computationally expensive end-to-end deep learning frameworks. By leveraging Linear Discriminant Analysis (LDA) for dimensionality reduction and Support Vector Classifier (SVC) for classification, our system enhances both accuracy and inference speed. The results confirm that lightweight architectures like MobileNetV2, when paired with appropriate classifiers and optimization techniques, can deliver state-of-the-art performance without excessive computational demands. Future research could explore self-supervised learning and transformer-based architectures, which have shown promise in medical imaging tasks. Additionally, incorporating multi-modal data sources, such as clinical history and radiological findings, could improve diagnostic accuracy. Expanding the dataset with more diverse and real-world stroke cases will further validate the model’s generalizability. Finally, developing a real-time deployment system with interpretability features will enhance clinical adoption, aiding medical professionals in fast and accurate stroke diagnosis.
Data Availability